A 2kg block is released 4m from a massless spring with a force constant K=100N/m that is aong a plane inlined at 30°. If the coefficient of kinetic friction between the plane and the block is 0.2, find the maximum compression.

1
Expert's answer
2021-12-09T09:15:00-0500

Explanations & Calculations


  • Refer to the sketch attached.
  • Also, refer to the algebraic solution which you can substitute values into to get the required answer.


  • Consider conservation of mechanical energy from location A to C. Here the energy is measured with respect to C at which we consider the level of zero potential energy.

"\qquad\qquad
\begin{aligned}
\small mgh&=\small \frac{1}{2}kx^2+f(d+x)\\
\small mgh&=\small \frac{1}{2}kx^2+\mu_kR(d+x)\\
\small mg([d+x]\sin\theta)&=\small \frac{1}{2}kx^2+\mu_kmg\cos\theta.(d+x)\\
\end{aligned}"


  • Here R is the normal force on the block from the incline, and it is equal to the component of the weight of the box that is perpendicular to the incline.


  • Here "\small d=4m" , "\small \theta =30^0," "\small k=100\,Nm^{-1}", "\small \mu_k=0.2" and "\small x =" the compression.
  • You can substitute and get the answer, here you will get a quadratic function of x you can solve it for x.